# A Genealogy of Multi-Sensor Foundation Models in Remote Sensing

### v1

@@ 7.4. Directions for Framework Development
Many of the foundation models covered in this manuscript leverage a variety of different objectives that force the encoder to learn accurate representations of the data and test on downstream generalizability. Most of these models use a single pretext task training objective in the hopes of generating a general representation. In other domains, such as NLP, multiple pretext tasks are optimized simultaneously as a key part of learning generalizable embeddings (Koroteev, 2021). RS foundation models could profit from a similar strategy.
@@
MMEarth (Nedungadi et al., 2024) utilizes multi-task training, as the model uses a variety of downstream decoders for different modalities that each backpropagate loss to the same encoder. FG-MAE (Wang et al., 2023a) takes this a step further by tasking decoders to not just reconstruct a different modality, but to reconstruct specific RS feature maps pertaining to the masked training image. Future research can investigate models that marry these MAE approaches with contrastive learning. Training a MAE encoder per modality, then performing contrastive loss between the resulting encoders could provide a powerful boost in representation learning, as both objectives learn invariant features in very different ways. Current foundation models that take text in as a modality inherently use this approach somewhat, as they rely on pre-trained NLP transformers, which have performed Masked Language Modeling training already for their text encoders (Koroteev, 2021). CROMA (Fuller et al., 2023) begins to explore this possibility in the RS domain by contrasting two independently masked images, but this is explored mainly for performance reasons, and the MAE objective is on multi-sensed image embeddings as opposed to MAE for each sensor. This approach may scale poorly with the addition of more sensors due to input size. Similarly, the nature of a MAE task shared with multiple modalities seems to align these sensors more closely, rendering it somewhat redundant to the contrastive task already being executed.
@@
Most multi-modal models in the RS domain only consider two to three sensors, but the sheer number of modes available means that it would be beneficial to have a model that is capable of scaling effectively to learn representations that capture a dynamic number of modes. MMEarth’s innovation of having a decoder per mode is promising in that regard, but it runs into the problem of only learning embeddings for one sensor. While contrastive learning can theoretically be applied across a variety of modes, the cost of training a separate encoder per sensor could quickly result in resource challenges. Ultimately, a multi-modal framework that is capable of learning representations for multiple sensors without explicitly requiring that every mode has its own encoder would be ideal, but careful consideration is required to develop a framework that is both flexible and robust enough. DOFA’s (Xiong et al., 2024) hypernetwork-inspired framework is promising in that regard due to its ability to handle a variety of different sensor inputs in pre-training, but it fails to take into account whether a wavelength is being actively sensed or passively sensed, which affects the nature of the imagery. Similarly, the various polarization of SAR acquisitions, even by the same sensor, can create further complications that go beyond wavelength: The C-band Sentinel-1 SAR images land areas in VH and VV mode, whereas it switches to HH and HV polarizations when over oceans (this is driven by the scientific and operational users need to map sea ice, which is better detectable in these modes). DOFA also does not demonstrably test what we consider to be this framework’s most interesting potential capability: the capacity to operate effectively on downstream tasks that rely on a sensor not found in the training dataset.
@@
SeCo (Mañas et al., 2021) and GASSL (Ayush et al., 2021) demonstrated the value of utilizing temporal data in training, but cautioned that the learned representations were agnostic to seasonal changes. The seasonal contrastive objective could be helpful as an additional task, when taken into account with other pre-training tasks that take seasonal change into account. Ultimately, multi-task training on complementary tasks, such as this particular example, is somewhat costly, but it results in representations that learn the unique features inherent to each task while avoiding the problem of overfitting. Aside from the SeCo objective, SatMAE’s (Cong et al., 2023) breakthrough in adding temporal information to the positional encoding of images could serve as a useful way to introduce temporal information into existing methods. However, caution should be taken to ensure that such encoders are robust enough to not be overly reliant on temporal information. ScaleMAE’s (Reed et al., 2023) novel approach of normalizing positional embeddings to account for image scale also seems to be a methodology that could be easily applied to other models within the field, presuming those models utilize a ViT framework.
@@ 8. Conclusion
In this paper we presented on a variety of SSL approaches currently utilized in RS foundation models and traced these approaches back to their origins in the CV field to develop a better understanding of the current swathe of RS foundation models and potential paths for future development. We examined:
@@ Acknowledgements.
- Alsweiss et al. (2015) Suleiman O. Alsweiss, Zorana Jelenak, Paul S. Chang, Jun Dong Park, and Patrick Meyers. 2015. Inter-calibration Results of the Advanced Microwave Scanning Radiometer-2 Over Ocean. IEEE Journal of Selected Topics in Applied Earth Observations and Remote Sensing 8, 9 (Sept. 2015), 4230–4238. doi:10.1109/JSTARS.2014.2330980 Conference Name: IEEE Journal of Selected Topics in Applied Earth Observations and Remote Sensing.
- Anderson and Marchisio (2012) Neal T. Anderson and Giovanni B. Marchisio. 2012. WorldView-2 and the evolution of the DigitalGlobe remote sensing satellite constellation: introductory paper for the special session on WorldView-2. In Algorithms and Technologies for Multispectral, Hyperspectral, and Ultraspectral Imagery XVIII, Vol. 8390. SPIE, 166–180. doi:10.1117/12.919756
@@
- Bailey et al. (2018) Michael Bailey, Rachel Cao, Theresa Kuchler, Johannes Stroebel, and Arlene Wong. 2018. Social connectedness: Measurement, determinants, and effects. Journal of Economic Perspectives 32, 3 (2018), 259–280.
- Barowski et al. (2019) Jan Barowski, Jochen Jebramcik, Jonas Wagner, Nils Pohl, and Ilona Rolfes. 2019. Spatial Identification of Dielectric Properties using Synthetic Aperture Radar. In 2019 IEEE MTT-S International Microwave Workshop Series on Advanced Materials and Processes for RF and THz Applications (IMWS-AMP). 139–141. doi:10.1109/IMWS-AMP.2019.8880121
- Brown et al. (2022) Christopher F. Brown, Steven P. Brumby, Brookie Guzder-Williams, Tanya Birch, Samantha Brooks Hyde, Joseph Mazzariello, Wanda Czerwinski, Valerie J. Pasquarella, Robert Haertel, Simon Ilyushchenko, Kurt Schwehr, Mikaela Weisse, Fred Stolle, Craig Hanson, Oliver Guinan, Rebecca Moore, and Alexander M. Tait. 2022. Dynamic World, Near real-time global 10 m land use land cover mapping. Scientific Data 9, 1 (June 2022), 251. doi:10.1038/s41597-022-01307-4
- Brown and Porcello (1969) William M. Brown and Leonard J. Porcello. 1969. An introduction to synthetic-aperture radar. IEEE Spectrum 6, 9 (Sept. 1969), 52–62. doi:10.1109/MSPEC.1969.5213674 Conference Name: IEEE Spectrum.
- Canny (1986) John Canny. 1986. A Computational Approach to Edge Detection. IEEE Transactions on Pattern Analysis and Machine Intelligence PAMI-8, 6 (Nov. 1986), 679–698. doi:10.1109/TPAMI.1986.4767851 Conference Name: IEEE Transactions on Pattern Analysis and Machine Intelligence.
- Caron et al. (2021) Mathilde Caron, Hugo Touvron, Ishan Misra, Hervé Jégou, Julien Mairal, Piotr Bojanowski, and Armand Joulin. 2021. Emerging Properties in Self-Supervised Vision Transformers. doi:10.48550/arXiv.2104.14294 arXiv:2104.14294 [cs].
- Castelletti et al. (2021) Davide Castelletti, Gordon Farquharson, Craig Stringham, Michael Duersch, and Duncan Eddy. 2021. Capella Space First Operational SAR Satellite. In 2021 IEEE International Geoscience and Remote Sensing Symposium IGARSS. 1483–1486. doi:10.1109/IGARSS47720.2021.9554100 ISSN: 2153-7003.
@@
- Chauhan et al. (2024) Vinod Kumar Chauhan, Jiandong Zhou, Ping Lu, Soheila Molaei, and David A. Clifton. 2024. A brief review of hypernetworks in deep learning. Artificial Intelligence Review 57, 9 (Sept. 2024), 1–29. doi:10.1007/s10462-024-10862-8 Company: Springer Distributor: Springer Institution: Springer Label: Springer Number: 9 Publisher: Springer Netherlands.
@@
- Chen et al. (2024) Yile Chen, Weiming Huang, Kaiqi Zhao, Yue Jiang, and Gao Cong. 2024. Self-supervised Learning for Geospatial AI: A Survey. doi:10.48550/arXiv.2408.12133 arXiv:2408.12133 [cs].
@@
- Chen et al. (2023) Y. Chen, X. S. Yu, and K. Qin. 2023. MMCPP: A MULTI-MODAL CONTRASTIVE PRE-TRAINING MODEL FOR PLACE REPRESENTATION BASED ON THE SPATIO-TEMPORAL FRAMEWORK. ISPRS Annals of the Photogrammetry, Remote Sensing and Spatial Information Sciences X-1-W1-2023 (Dec. 2023), 303–310. doi:10.5194/isprs-annals-X-1-W1-2023-303-2023 Conference Name: ISPRS Geospatial Week 2023 - 2–7 September 2023, Cairo, Egypt Publisher: Copernicus GmbH.
@@
- Cheng et al. (2017) Gong Cheng, Junwei Han, and Xiaoqiang Lu. 2017. Remote Sensing Image Scene Classification: Benchmark and State of the Art. Proc. IEEE 105, 10 (Oct. 2017), 1865–1883. doi:10.1109/JPROC.2017.2675998 Conference Name: Proceedings of the IEEE.
@@
- Christie et al. (2018) Gordon Christie, Neil Fendley, James Wilson, and Ryan Mukherjee. 2018. Functional Map of the World. 6172–6180. https://openaccess.thecvf.com/content_cvpr_2018/html/Christie_Functional_Map_of_CVPR_2018_paper.html
@@
- Cillero Castro et al. (2020) Carmen Cillero Castro, Jose Antonio Domínguez Gómez, Jordi Delgado Martín, Boris Alejandro Hinojo Sánchez, Jose Luis Cereijo Arango, Federico Andrés Cheda Tuya, and Ramon Díaz-Varela. 2020. An UAV and Satellite Multispectral Data Approach to Monitor Water Quality in Small Reservoirs. Remote Sensing 12, 9 (Jan. 2020), 1514. doi:10.3390/rs12091514 Number: 9 Publisher: Multidisciplinary Digital Publishing Institute.
@@
- Claverie et al. (2018) Martin Claverie, Junchang Ju, Jeffrey G. Masek, Jennifer L. Dungan, Eric F. Vermote, Jean-Claude Roger, Sergii V. Skakun, and Christopher Justice. 2018. The Harmonized Landsat and Sentinel-2 surface reflectance data set. Remote Sensing of Environment 219 (Dec. 2018), 145–161. doi:10.1016/j.rse.2018.09.002
@@
- Cong et al. (2023) Yezhen Cong, Samar Khanna, Chenlin Meng, Patrick Liu, Erik Rozi, Yutong He, Marshall Burke, David B. Lobell, and Stefano Ermon. 2023. SatMAE: Pre-training Transformers for Temporal and Multi-Spectral Satellite Imagery. doi:10.48550/arXiv.2207.08051 arXiv:2207.08051 [cs].
@@
- Dalal and Triggs (2005) N. Dalal and B. Triggs. 2005. Histograms of oriented gradients for human detection. In 2005 IEEE Computer Society Conference on Computer Vision and Pattern Recognition (CVPR’05), Vol. 1. 886–893 vol. 1. doi:10.1109/CVPR.2005.177 ISSN: 1063-6919.
@@
- Daudt et al. (2018) Rodrigo Caye Daudt, Bertr Le Saux, Alexandre Boulch, and Yann Gousseau. 2018. Urban Change Detection for Multispectral Earth Observation Using Convolutional Neural Networks. In IGARSS 2018 - 2018 IEEE International Geoscience and Remote Sensing Symposium. 2115–2118. doi:10.1109/IGARSS.2018.8518015 ISSN: 2153-7003.
@@
- Dehghani et al. (2023) Mostafa Dehghani, Josip Djolonga, Basil Mustafa, Piotr Padlewski, Jonathan Heek, Justin Gilmer, Andreas Steiner, Mathilde Caron, Robert Geirhos, Ibrahim Alabdulmohsin, Rodolphe Jenatton, Lucas Beyer, Michael Tschannen, Anurag Arnab, Xiao Wang, Carlos Riquelme, Matthias Minderer, Joan Puigcerver, Utku Evci, Manoj Kumar, Sjoerd van Steenkiste, Gamaleldin F. Elsayed, Aravindh Mahendran, Fisher Yu, Avital Oliver, Fantine Huot, Jasmijn Bastings, Mark Patrick Collier, Alexey Gritsenko, Vighnesh Birodkar, Cristina Vasconcelos, Yi Tay, Thomas Mensink, Alexander Kolesnikov, Filip Pavetić, Dustin Tran, Thomas Kipf, Mario Lučić, Xiaohua Zhai, Daniel Keysers, Jeremiah Harmsen, and Neil Houlsby. 2023. Scaling Vision Transformers to 22 Billion Parameters. doi:10.48550/arXiv.2302.05442 arXiv:2302.05442 [cs].
@@
- Entekhabi et al. (2010) Dara Entekhabi, Eni G. Njoku, Peggy E. O’Neill, Kent H. Kellogg, Wade T. Crow, Wendy N. Edelstein, Jared K. Entin, Shawn D. Goodman, Thomas J. Jackson, Joel Johnson, John Kimball, Jeffrey R. Piepmeier, Randal D. Koster, Neil Martin, Kyle C. McDonald, Mahta Moghaddam, Susan Moran, Rolf Reichle, J. C. Shi, Michael W. Spencer, Samuel W. Thurman, Leung Tsang, and Jakob Van Zyl. 2010. The Soil Moisture Active Passive (SMAP) Mission. Proc. IEEE 98, 5 (May 2010), 704–716. doi:10.1109/JPROC.2010.2043918 Conference Name: Proceedings of the IEEE.

### v2

@@ 10.4. Directions for Framework Development
Many of the foundation models covered in this manuscript leverage a variety of different objectives that force the encoder to learn accurate representations of the data and test on downstream generalizability. Many of these models use a single pretext task training objective in the hopes of generating a general representation. In other domains, such as NLP, multiple pretext tasks are optimized simultaneously as a key part of learning generalizable embeddings (Koroteev, 2021). RS foundation models could profit from a similar strategy.
@@
MMEarth (Nedungadi et al., 2024) utilizes multi-task training, as the model uses a variety of downstream decoders for different modalities that each backpropagate loss to the same encoder. FG-MAE (Wang et al., 2023b) takes this a step further by tasking decoders to not just reconstruct a different modality, but to reconstruct specific RS feature maps pertaining to the masked training image. While we have discussed some approaches designed to leverage contrastive loss between modes that also include MAE within a mode, this remains an under-explored avenue. CROMA (Fuller et al., 2023) begins to explore this possibility in the RS domain by contrasting two independently masked images, but this is explored mainly for performance reasons, and the MAE objective is on multi-sensed image embeddings as opposed to MAE for each sensor. This approach may scale poorly with the addition of more sensors due to input size. Similarly, the nature of a MAE task shared with multiple modalities seems to align these sensors more closely, rendering it somewhat redundant to the contrastive task already being executed. DeCUR (Wang et al., 2024) provides an interesting alternative to such approaches. It leverages the same redundancy reduction task across modes and within modes, but the choice to set aside roughly 20% of the latent space dimensions for learning modality specific features seems to be a framework choice that other foundation models could benefit from. Similarly, the regularization term in VICReg (Bardes et al., 2022) that enforces a degree of standard deviation within each dimension is worth exploring as an avenue to prevent embedding collapse.
@@
Most multi-modal models in the RS domain only consider two to three sensors, but the sheer number of modes available means that it would be beneficial to have a model that is capable of scaling effectively to learn representations that capture a dynamic number of modes. MMEarth’s innovation of having a decoder per mode is promising in that regard, but it runs into the problem of only learning embeddings for one sensor. While contrastive learning can theoretically be applied across a variety of modes, the cost of training a separate encoder per sensor could quickly result in resource challenges. For example, OmniSat (Astruc et al., 2024) leverages MAE per mode and contrastive loss across three modes, but the contrastive framework would scale poorly with the addition of more modalities. Ultimately, a multi-modal framework that is capable of learning representations for multiple sensors without explicitly requiring that every mode has its own encoder would be ideal, but careful consideration is required to develop a framework that is both flexible and robust enough. One potential option would be following an ImageBind (Girdhar et al., 2023) approach, where one modality is the ’binding’ modality that all other modalities are contrasted to. Treating spatiotemporal location as a binding modality for any remotely sensed data could be an interesting avenue to pursue. Such a framework would allow incorporation of pre-trained, sensor-specific encoders. The number of comparisons made would scale linearly with the number of additional modalities, as opposed to exponentially.
DOFA’s (Xiong et al., 2024) hypernetwork-inspired framework is promising in that regard due to its ability to handle a variety of different sensor inputs in pre-training, but it fails to take into account whether a wavelength is being actively sensed or passively sensed, which affects the nature of the imagery. Similarly, the various polarization of SAR acquisitions, even by the same sensor, can create further complications that go beyond wavelength: The C-band Sentinel-1 SAR images land areas in VH and VV mode, whereas it switches to HH and HV polarizations when over oceans (this is driven by the scientific and operational users’ need to map sea ice, which is better detectable in these modes). DOFA also does not demonstrably test what we consider to be this framework’s most interesting potential capability: the capacity to operate effectively on downstream tasks that rely on a sensor not found in the training dataset.
@@
SeCo (Mañas et al., 2021) and GASSL (Ayush et al., 2021) demonstrated the value of utilizing temporal data in training, but cautioned that the learned representations were agnostic to seasonal changes. The seasonal contrastive objective could be helpful as an additional task, when taken into account with other pre-training tasks that take seasonal change into account. Ultimately, multi-task training on complementary tasks, such as this particular example, is somewhat costly, but it results in representations that learn the unique features inherent to each task while avoiding the problem of overfitting. Aside from the SeCo objective, SatMAE’s (Cong et al., 2023) breakthrough in adding temporal information to the positional encoding of images could serve as a useful way to introduce temporal information into existing methods. However, caution should be taken to ensure that such encoders are robust enough to not be overly reliant on temporal information. ScaleMAE’s (Reed et al., 2023) novel approach of normalizing positional embeddings to account for image scale also seems to be a methodology that could be easily applied to other models within the field, presuming those models utilize a ViT framework. Similarly, Cross-Scale MAE’s (Tang et al., 2024) focus on random crop and resize augmentations as a way to learn spatial resolution could be applied to other contrastive and distillation based frameworks.
@@ 11. Conclusion
In this paper we presented on a variety of SSL approaches currently utilized in RS foundation models. We traced these approaches back to their origins in the CV field to develop a better understanding of the current swathe of RS foundation models and potential paths for future development.
We examined:
@@ Acknowledgements.
- Alsweiss et al. (2015) Suleiman O. Alsweiss, Zorana Jelenak, Paul S. Chang, Jun Dong Park, and Patrick Meyers. 2015. Inter-calibration Results of the Advanced Microwave Scanning Radiometer-2 Over Ocean. IEEE Journal of Selected Topics in Applied Earth Observations and Remote Sensing 8, 9 (Sept. 2015), 4230–4238. https://doi.org/10.1109/JSTARS.2014.2330980 Conference Name: IEEE Journal of Selected Topics in Applied Earth Observations and Remote Sensing.
- Anderson and Marchisio (2012) Neal T. Anderson and Giovanni B. Marchisio. 2012. WorldView-2 and the evolution of the DigitalGlobe remote sensing satellite constellation: introductory paper for the special session on WorldView-2. In Algorithms and Technologies for Multispectral, Hyperspectral, and Ultraspectral Imagery XVIII, Vol. 8390. SPIE, 166–180. https://doi.org/10.1117/12.919756
- Anisuzzaman et al. (2025) D. M. Anisuzzaman, Jeffrey G. Malins, Paul A. Friedman, and Zachi I. Attia. 2025. Fine-Tuning Large Language Models for Specialized Use Cases. Mayo Clinic Proceedings: Digital Health 3, 1 (March 2025), 100184. https://doi.org/10.1016/j.mcpdig.2024.11.005
- Assran et al. (2023) Mahmoud Assran, Quentin Duval, Ishan Misra, Piotr Bojanowski, Pascal Vincent, Michael Rabbat, Yann LeCun, and Nicolas Ballas. 2023. Self-Supervised Learning From Images With a Joint-Embedding Predictive Architecture. 15619–15629. https://openaccess.thecvf.com/content/CVPR2023/html/Assran_Self-Supervised_Learning_From_Images_With_a_Joint-Embedding_Predictive_Architecture_CVPR_2023_paper.html
- Astruc et al. (2024) Guillaume Astruc, Nicolas Gonthier, Clement Mallet, and Loic Landrieu. 2024. OmniSat: Self-Supervised Modality Fusion for Earth Observation. https://doi.org/10.48550/arXiv.2404.08351 arXiv:2404.08351 [cs].
@@
- Baier et al. (2020) Gerald Baier, Antonin Deschemps, Michael Schmitt, and Naoto Yokoya. 2020. GeoNRW. https://doi.org/10.21227/S5XQ-B822
- Bailey et al. (2018) Michael Bailey, Rachel Cao, Theresa Kuchler, Johannes Stroebel, and Arlene Wong. 2018. Social Connectedness: Measurement, Determinants, and Effects. Journal of Economic Perspectives 32, 3 (Aug. 2018), 259–280. https://doi.org/10.1257/jep.32.3.259
- Bao et al. (2022) Hangbo Bao, Li Dong, Songhao Piao, and Furu Wei. 2022. BEiT: BERT Pre-Training of Image Transformers. https://doi.org/10.48550/arXiv.2106.08254 arXiv:2106.08254 [cs].
- Bardes et al. (2022) Adrien Bardes, Jean Ponce, and Yann LeCun. 2022. VICReg: Variance-Invariance-Covariance Regularization for Self-Supervised Learning. https://doi.org/10.48550/arXiv.2105.04906 arXiv:2105.04906 [cs].
- Barowski et al. (2019) Jan Barowski, Jochen Jebramcik, Jonas Wagner, Nils Pohl, and Ilona Rolfes. 2019. Spatial Identification of Dielectric Properties using Synthetic Aperture Radar. In 2019 IEEE MTT-S International Microwave Workshop Series on Advanced Materials and Processes for RF and THz Applications (IMWS-AMP). 139–141. https://doi.org/10.1109/IMWS-AMP.2019.8880121
- Bonafilia et al. (2020) Derrick Bonafilia, Beth Tellman, Tyler Anderson, and Erica Issenberg. 2020. Sen1Floods11: A Georeferenced Dataset to Train and Test Deep Learning Flood Algorithms for Sentinel-1. 210–211. https://openaccess.thecvf.com/content_CVPRW_2020/html/w11/Bonafilia_Sen1Floods11_A_Georeferenced_Dataset_to_Train_and_Test_Deep_Learning_CVPRW_2020_paper.html
- Bountos et al. (2025) Nikolaos Ioannis Bountos, Arthur Ouaknine, Ioannis Papoutsis, and David Rolnick. 2025. FoMo: Multi-Modal, Multi-Scale and Multi-Task Remote Sensing Foundation Models for Forest Monitoring. Proceedings of the AAAI Conference on Artificial Intelligence 39, 27 (April 2025), 27858–27868. https://doi.org/10.1609/aaai.v39i27.35002
- Brown et al. (2022) Christopher F. Brown, Steven P. Brumby, Brookie Guzder-Williams, Tanya Birch, Samantha Brooks Hyde, Joseph Mazzariello, Wanda Czerwinski, Valerie J. Pasquarella, Robert Haertel, Simon Ilyushchenko, Kurt Schwehr, Mikaela Weisse, Fred Stolle, Craig Hanson, Oliver Guinan, Rebecca Moore, and Alexander M. Tait. 2022. Dynamic World, Near real-time global 10 m land use land cover mapping. Scientific Data 9, 1 (June 2022), 251. https://doi.org/10.1038/s41597-022-01307-4
- Brown and Porcello (1969) William M. Brown and Leonard J. Porcello. 1969. An introduction to synthetic-aperture radar. IEEE Spectrum 6, 9 (Sept. 1969), 52–62. https://doi.org/10.1109/MSPEC.1969.5213674 Conference Name: IEEE Spectrum.
- Burt and Adelson (1983) P. Burt and E. Adelson. 1983. The Laplacian Pyramid as a Compact Image Code. IEEE Transactions on Communications 31, 4 (April 1983), 532–540. https://doi.org/10.1109/TCOM.1983.1095851 Conference Name: IEEE Transactions on Communications.
- Canny (1986) John Canny. 1986. A Computational Approach to Edge Detection. IEEE Transactions on Pattern Analysis and Machine Intelligence PAMI-8, 6 (Nov. 1986), 679–698. https://doi.org/10.1109/TPAMI.1986.4767851 Conference Name: IEEE Transactions on Pattern Analysis and Machine Intelligence.
- Caron et al. (2021) Mathilde Caron, Hugo Touvron, Ishan Misra, Hervé Jégou, Julien Mairal, Piotr Bojanowski, and Armand Joulin. 2021. Emerging Properties in Self-Supervised Vision Transformers. https://doi.org/10.48550/arXiv.2104.14294 arXiv:2104.14294 [cs].
- Castelletti et al. (2021) Davide Castelletti, Gordon Farquharson, Craig Stringham, Michael Duersch, and Duncan Eddy. 2021. Capella Space First Operational SAR Satellite. In 2021 IEEE International Geoscience and Remote Sensing Symposium IGARSS. 1483–1486. https://doi.org/10.1109/IGARSS47720.2021.9554100 ISSN: 2153-7003.
@@
- Chauhan et al. (2024) Vinod Kumar Chauhan, Jiandong Zhou, Ping Lu, Soheila Molaei, and David A. Clifton. 2024. A brief review of hypernetworks in deep learning. Artificial Intelligence Review 57, 9 (Sept. 2024), 1–29. https://doi.org/10.1007/s10462-024-10862-8 Company: Springer Distributor: Springer Institution: Springer Label: Springer Number: 9 Publisher: Springer Netherlands.
@@
- Chen et al. (2024) Yile Chen, Weiming Huang, Kaiqi Zhao, Yue Jiang, and Gao Cong. 2024. Self-supervised Learning for Geospatial AI: A Survey. https://doi.org/10.48550/arXiv.2408.12133 arXiv:2408.12133 [cs].
@@
- Chen et al. (2023) Y. Chen, X. S. Yu, and K. Qin. 2023. MMCPP: A MULTI-MODAL CONTRASTIVE PRE-TRAINING MODEL FOR PLACE REPRESENTATION BASED ON THE SPATIO-TEMPORAL FRAMEWORK. ISPRS Annals of the Photogrammetry, Remote Sensing and Spatial Information Sciences X-1-W1-2023 (Dec. 2023), 303–310. https://doi.org/10.5194/isprs-annals-X-1-W1-2023-303-2023 Conference Name: ISPRS Geospatial Week 2023 - 2–7 September 2023, Cairo, Egypt Publisher: Copernicus GmbH.
- Chen et al. (2022) Zhengyang Chen, Yao Qian, Bing Han, Yanmin Qian, and Michael Zeng. 2022. A comprehensive study on self-supervised distillation for speaker representation learning. https://doi.org/10.48550/arXiv.2210.15936 arXiv:2210.15936 [cs].
@@
- Cheng et al. (2017) Gong Cheng, Junwei Han, and Xiaoqiang Lu. 2017. Remote Sensing Image Scene Classification: Benchmark and State of the Art. Proc. IEEE 105, 10 (Oct. 2017), 1865–1883. https://doi.org/10.1109/JPROC.2017.2675998 Conference Name: Proceedings of the IEEE.
@@
- Christie et al. (2018) Gordon Christie, Neil Fendley, James Wilson, and Ryan Mukherjee. 2018. Functional Map of the World. https://doi.org/10.48550/arXiv.1711.07846 arXiv:1711.07846 [cs].
@@
- Cillero Castro et al. (2020) Carmen Cillero Castro, Jose Antonio Domínguez Gómez, Jordi Delgado Martín, Boris Alejandro Hinojo Sánchez, Jose Luis Cereijo Arango, Federico Andrés Cheda Tuya, and Ramon Díaz-Varela. 2020. An UAV and Satellite Multispectral Data Approach to Monitor Water Quality in Small Reservoirs. Remote Sensing 12, 9 (Jan. 2020), 1514. https://doi.org/10.3390/rs12091514 Number: 9 Publisher: Multidisciplinary Digital Publishing Institute.
@@
- Claverie et al. (2018) Martin Claverie, Junchang Ju, Jeffrey G. Masek, Jennifer L. Dungan, Eric F. Vermote, Jean-Claude Roger, Sergii V. Skakun, and Christopher Justice. 2018. The Harmonized Landsat and Sentinel-2 surface reflectance data set. Remote Sensing of Environment 219 (Dec. 2018), 145–161. https://doi.org/10.1016/j.rse.2018.09.002
@@
- Cong et al. (2023) Yezhen Cong, Samar Khanna, Chenlin Meng, Patrick Liu, Erik Rozi, Yutong He, Marshall Burke, David B. Lobell, and Stefano Ermon. 2023. SatMAE: Pre-training Transformers for Temporal and Multi-Spectral Satellite Imagery. https://doi.org/10.48550/arXiv.2207.08051 arXiv:2207.08051 [cs].
@@
- Dalal and Triggs (2005) N. Dalal and B. Triggs. 2005. Histograms of oriented gradients for human detection. In 2005 IEEE Computer Society Conference on Computer Vision and Pattern Recognition (CVPR’05), Vol. 1. 886–893 vol. 1. https://doi.org/10.1109/CVPR.2005.177 ISSN: 1063-6919.
@@
- Daudt et al. (2018) Rodrigo Caye Daudt, Bertr Le Saux, Alexandre Boulch, and Yann Gousseau. 2018. Urban Change Detection for Multispectral Earth Observation Using Convolutional Neural Networks. In IGARSS 2018 - 2018 IEEE International Geoscience and Remote Sensing Symposium. 2115–2118. https://doi.org/10.1109/IGARSS.2018.8518015 ISSN: 2153-7003.
@@
- Dehghani et al. (2023) Mostafa Dehghani, Josip Djolonga, Basil Mustafa, Piotr Padlewski, Jonathan Heek, Justin Gilmer, Andreas Steiner, Mathilde Caron, Robert Geirhos, Ibrahim Alabdulmohsin, Rodolphe Jenatton, Lucas Beyer, Michael Tschannen, Anurag Arnab, Xiao Wang, Carlos Riquelme, Matthias Minderer, Joan Puigcerver, Utku Evci, Manoj Kumar, Sjoerd van Steenkiste, Gamaleldin F. Elsayed, Aravindh Mahendran, Fisher Yu, Avital Oliver, Fantine Huot, Jasmijn Bastings, Mark Patrick Collier, Alexey Gritsenko, Vighnesh Birodkar, Cristina Vasconcelos, Yi Tay, Thomas Mensink, Alexander Kolesnikov, Filip Pavetić, Dustin Tran, Thomas Kipf, Mario Lučić, Xiaohua Zhai, Daniel Keysers, Jeremiah Harmsen, and Neil Houlsby. 2023. Scaling Vision Transformers to 22 Billion Parameters. https://doi.org/10.48550/arXiv.2302.05442 arXiv:2302.05442 [cs].
- Devlin et al. (2019) Jacob Devlin, Ming-Wei Chang, Kenton Lee, and Kristina Toutanova. 2019. BERT: Pre-training of Deep Bidirectional Transformers for Language Understanding. https://doi.org/10.48550/arXiv.1810.04805 arXiv:1810.04805 [cs].
- Dosovitskiy et al. (2021) Alexey Dosovitskiy, Lucas Beyer, Alexander Kolesnikov, Dirk Weissenborn, Xiaohua Zhai, Thomas Unterthiner, Mostafa Dehghani, Matthias Minderer, Georg Heigold, Sylvain Gelly, Jakob Uszkoreit, and Neil Houlsby. 2021. An Image is Worth 16x16 Words: Transformers for Image Recognition at Scale. https://doi.org/10.48550/arXiv.2010.11929 arXiv:2010.11929 [cs].
@@
- Entekhabi et al. (2010) Dara Entekhabi, Eni G. Njoku, Peggy E. O’Neill, Kent H. Kellogg, Wade T. Crow, Wendy N. Edelstein, Jared K. Entin, Shawn D. Goodman, Thomas J. Jackson, Joel Johnson, John Kimball, Jeffrey R. Piepmeier, Randal D. Koster, Neil Martin, Kyle C. McDonald, Mahta Moghaddam, Susan Moran, Rolf Reichle, J. C. Shi, Michael W. Spencer, Samuel W. Thurman, Leung Tsang, and Jakob Van Zyl. 2010. The Soil Moisture Active Passive (SMAP) Mission. Proc. IEEE 98, 5 (May 2010), 704–716. https://doi.org/10.1109/JPROC.2010.2043918 Conference Name: Proceedings of the IEEE.
- Fedus et al. (2022) William Fedus, Barret Zoph, and Noam Shazeer. 2022. Switch Transformers: Scaling to Trillion Parameter Models with Simple and Efficient Sparsity. https://doi.org/10.48550/arXiv.2101.03961 arXiv:2101.03961 [cs].